OpenStudy (anonymous):

what is 3loga+1/2logb use logarith properties to rewrite as a single logarithm

OpenStudy (anonymous):

= loga^3 + logb^(1/2) = loga^3b^(1/2)

OpenStudy (anonymous):

\[\log a ^{3}\sqrt{b}\]

OpenStudy (anonymous):

\[ \log( {a^{3} \times \sqrt{b} })\] yeah they're right